# Twisting Lids Off with Two Hands

### v1

@@ IV-A Real-world System
IV-B Bottle Models
Simulated Bottles. A central challenge in simulating the lid-twisting task is how to properly model friction between the bottle body and the lid, particularly static friction. Simulating this type of physical force has been a long-standing problem in graphics [33]. We design a simple modeling approximation that strikes a balance between fidelity and speed during physical simulation. As a result, a successful policy can be learned in simulation within a reasonable amount of wall-clock time while being realistic enough for transfer to the real world. We illustrate our bottle-like object model in Figure 1. Our design features a special Brake Link that constantly presses against the bottle lid (through a prismatic joint). This artificially generates frictional forces between the bottle body (Base Link) and the lid (Lid Link), preventing relative rotation between them—similar to a bottle with its lid screwed on.
Real-World Articulated Bottles. To easily vary the shapes and sizes of our objects and encourage reproducibility, we design a set of articulated bottle-like objects in CAD software and 3D print them in a range of different colors and materials. The lids can be rotated infinitely, simulating an ideal revolute joint. We show examples of our custom bottles in Figure 4 (middle). We also collect household bottle-like objects with varying physical properties to demonstrate the generalizability of our method, as depicted in Figure 4 (bottom).
V Learning to Twist Lids
Bimanual in-hand dexterous manipulation involves highly complex hand-object contacts, and remains challenging to solve with traditional methods. In this work, we address the control challenge through RL. We formulate our control problem as a partially observable Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{P})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{O}$ is the observation space, $\mathcal{R}$ is the reward function, and $\mathcal{P}$ is the environment dynamics. The control policy generates an action $a_{t}\in\mathcal{A}$ to maximize the cumulative reward given an observation $o_{t}\in\mathcal{O}$.
V-A Observation and Action
Action Space. We use an impedance PD controller to drive the robot hand. The control policy produces a relative target joint position as the action $a_{t}$, which is added to the current target joint position $q_{t}$ to produce the next target: $\tilde{q}_{t+1}=\tilde{q}_{t}+\eta{\rm EMA}(a_{t})$. $\eta$ is a scaling factor. Note that we smooth the action with its exponential moving average (EMA) to produce smooth motion. The next target position is sent to the PD controller to generate torque on each joint.
V-B Reward Design
Figure 3: Illustration of our reward design. Our task-specific reward contains three terms: finger contact reward (yellow arrows), twisting reward (white arrow), and pose reward (blue arrow). In particular, our keypoint-based finger reward is crucial for learning desired screwing behavior, which encourages the robot to minimize the distance between fingertips and reference contact keypoints.
The reward function plays a critical role in defining the robot’s behavior. The bimanual in-hand manipulation system has a higher action space dimension and much more complex contact patterns compared to single-hand scenarios. This magnifies the challenges inherent in the RL exploration process, underlining the need for more sophisticated reward designs. While one way to approach hard exploration problems is to add intrinsic rewards [4, 29], we propose a fine-grained, behavior-aware reward design for this task. Our reward function contains the following terms:
Twisting Reward. We define the twisting reward as
$r_{\mathrm{twisting}}=\Delta\theta=q^{t+1}_{bottle}-q^{t}_{bottle},$ (1)
which is the rotation angle of the lid during one-step execution.
Finger Contact Reward. We encourage natural object manipulation behavior by defining a set of keypoints positioned on the object to guide the fingers’ interactions (Figure 3). We define two set of keypoints $\mathbf{X}^{L}\in\mathbb{R}^{n\times 3}$ and $\mathbf{X}^{R}\in\mathbb{R}^{m\times 3}$ attached on the bottle base and lid respectively. Then, we define the finger contact reward as
$r_{\mathrm{contact}}=\sum_{i}\left[\frac{1}{1+\alpha d(\mathbf{X}^{L},\mathbf{F}^{L}_{i})}+\frac{1}{1+\alpha d(\mathbf{X}^{R},\mathbf{F}^{R}_{i})}\right],$ (2)
where $\mathbf{F}^{L}\in\mathbb{R}^{4\times 3}$ and $\mathbf{F}^{R}\in\mathbb{R}^{4\times 3}$ are the position of left and right fingertips, $\alpha$ is a scaling hyperparameter, and $d$ is a distance function defined as
$d(\mathbf{A},\mathbf{x})=\min_{i}\|\mathbf{A}_{i}-\mathbf{x}\|_{2}.$ (3)
Therefore, we encourage each fingertip to stay as close to one of the keypoints as possible. By choosing the keypoint properly around the bottle base and lid, we can elicit natural grasping and lid-twisting behavior.
Pose Reward. We also introduce a pose matching reward term to encourage the bottle main axis $\mathbf{x}_{axis}$ aligned with a predefined direction $\mathbf{v}$. This term is defined as
$r_{\mathrm{pose}}=-\arccos(\langle\mathbf{x}_{axis},\mathbf{v}\rangle).$ (4)
Other Regularizations. The above three reward terms specify the objective of our task. Besides these three rewards, we also introduce another few regularization terms as in previous works [55], including work penalty and action penalty to penalize large, jerky motions:
$r_{\mathrm{action}}=-\|a_{t}\|^{2},$ (5)
$r_{\mathrm{work}}=-\langle|\tau|,|\dot{q}_{t}|\rangle.$ (6)
The reward function is a weighted sum of the above terms:
$r=\alpha_{1}r_{\mathrm{contact}}+\alpha_{2}r_{\mathrm{twist}}+\alpha_{3}r_{\mathrm{pose}}+\alpha_{4}r_{\mathrm{action}}+\alpha_{5}r_{\mathrm{work}}.$ (7)
Reset Strategy. Due to the high dimensionality of our learning problem, there exist many possible ways how the robot interact with the object; among these, most modes lead to failures such as dropping the object beyond recovery. Exploring these modes rarely leads to meaningful learning progress. To circumvent this issue, we introduce several early termination criteria. Most importantly, we reset an episode if the robot hands fail to rotate the bottle into a desired pose for bimanual twisting within a short time limit. Additionally, we also reset when the bottle’s $z$-position is below a certain threshold, as the fingertips of the two hands can pinch the bottle at a low position without being able to reposition it into the palm.
V-C Domain Randomization
We apply a wide range of domain randomizations to ensure zero-shot sim-to-real transfer, including both physical and non-physical randomizations. Physical randomizations include the randomization of object friction, mass, and scale. We also apply random forces to the object to simulate the physical effects that are not implemented by the simulator. Non-physical randomizations model the noise in observation (e.g. joint position measurement and detected object positions) and action. A summary of our randomization attributes and parameters is shown in Table I.
TABLE I: Domain Randomization Setup.
Object: Mass (kg) [0.03, 0.1]
Object: Friction [0.5, 1.5]
Object: Shape ×$\times\mathcal{U}(0.95,1.05)$
Object: Initial Position (cm) $+\mathcal{U}(-0.02,0.02)$
Object: Initial $z$-orientation $+\mathcal{U}(-0.75,0.75)$
Hand: Friction [0.5, 1.5]
PD Controller: P Gain ×$\times\mathcal{U}(0.8,1.1)$
PD Controller: D Gain ×$\times\mathcal{U}(0.7,1.2)$
Random Force: Scale 2.0
Random Force: Probability 0.2
Random Force: Decay Coeff. and Interval 0.99 every 0.1s
Bottle Pos Observation: Noise 0.02
Joint Observation Noise. $+\mathcal{N}(0,0.4)$
Action Noise. $+\mathcal{N}(0,0.1)$
Frame Lag Probability 0.1
Action Lag Probability 0.1
V-D Training
We use the proximal policy optimization (PPO) algorithm to learn RL policies. We use an advantage clipping coefficient $\epsilon=0.2$; a horizon length of 16, with $\gamma=0.99$, and generalized advantage estimator (GAE) [46] coefficient $\tau=0.95$. The policy network is a three-layer MLP with ELU [13] activation, whose hidden layer is [256, 256, 128]. The policy network outputs a Gaussian distribution with a learnable state-independent standard deviation. The value network is also an MLP with ELU activation, whose hidden layer is [512, 512, 512]. We use an adaptive learning rate with KL threshold of 0.016 [19]. During training, we normalize the state input, value, and advantage. The gradient norm is set to 1.0 and the minibatch size is set to 8192. We use asymmetric observation [37] for the policy and value network, adding privileged information (randomized physical parameters, cube velocity and angular velocity, joint velocity and angular velocity) to the value network inputs. This privileged information is not accessible by the policy network.
Figure 4: Used bottles in our experiments. Top: Simulated bottles. Middle: Custom-made bottles (in-distribution except for the rightmost square bottle). Bottom: Household object bottles (out-of-distribution).
Figure 5: Training curves in different settings. Top: Results in single-object setup. Bottom: Results in multi-object setup. Left half: Comparisons of different reward setups. Right half: Ablations on the use of vision. The results are averaged on 5 seeds. The shaded area shows the standard deviation. The AD score is averaged by the total execution steps.
VI Simulated Experiments
We study the following questions in simulation:
Reward Design. How important is the keypoint-based reward for eliciting desired twisting behavior in bimanual manipulation? How effective is it compared to the other reward functions used in recent works?
Perception. How important is the visual information in solving this task? Is a sparse, keypoint representation enough for learning a policy that can handle multiple objects?
VI-A Setup
Object Set. In the simulated experiments, we utilize a collection of simulated cylindrical bottles with varying aspect ratios for both training and evaluation. Some samples are visualized in Figure 4. We consider two setups in simulation: 1) Multi-object setup, in which all the objects are used, and 2) Single-object setup, in which we only use a medium-sized bottle that represents the mean of the dataset.
Figure 6: Behavior of different reward functions. Top: Our full reward function achieves a stable grasp, as well as a smooth, natural, and human-like twisting motion. Middle: A naive gait constraint reward without any contact hints leads to erratic finger motion and unnatural grasps. Bottom: A reduced contact reward yields somewhat natural behavior, but the grasp is loose and ungainly compared to the full contact reward case.
Evaluation Metric. We introduce the following metrics for evaluating the performance:
Angular Displacement (AD) is the total number of degrees through which the lid has been twisted.
Time-to-Fail (TTF) is the period measured from the moment the bottle is held to the point when it either slips from the hand or becomes lodged.
Baselines. We compare our policy with the following baselines:
Policy without Vision. This is a neural network policy without visual information (bottle keypoints). We use this to evaluate the importance of vision.
Policy with Reduced Contact Reward. In training this policy, we reduce the intensity of our proposed finger contact reward. We use this to study the role of our contact reward in policy learning and shaping the policy’s behavior.
Policy with Gait Reward. In training this policy, we replace our contact reward with a gait constraint reward function similar to ones used for in-hand reorientation tasks (e.g. [40]). This baseline is only used for qualitative analysis.
VI-B Main Results
Reward Design. We first compare our approach with the reduced finger reward baseline (Figure 5). After decreasing the scale of finger contact reward, learned policies fail to master the desired lid-twisting skill and have low performance in general. We hypothesize that this is because the motion of lid-twisting requires a very specific pose pattern for holding the object; without explicitly encouraging such a pose pattern (e.g., via its contact modes), RL exploration becomes so hard that it is unsolvable within the available training time. We also observe a positive correlation between the intensity of finger contact reward and both 1) sample efficiency during learning and 2) performance of learned policies (as reflected by the AD score).
Vision vs No Vision. We also study the importance of vision modality in solving our considered task. Existing works show that certain rotation behavior can be achieved through implicit tactile sensing (joint proprioception data) [40], but it is unclear whether the same conclusion can be drawn for the bimanual lid-twisting task. Our empirical results show that, in both single and multi-object setups, the no-vision baseline performs substantially worse than our full method. This suggests that knowledge of the position of bottle keypoints is essential for successful lid-twisting.
Single Object vs Multi Object. We run RL training with two object settings: (1) using a single bottle-like object and (2) using multiple bottle-like objects with more variation in the ratio between the bottle base and lid. The two settings pose a trade-off between specialization and generalization: in the single-object scenario, the policy might learn successful behaviors more easily but find it harder to generalize to unseen objects, and vice versa. To our surprise, we observe that multi-object training yields slightly better performance compared to single-object training. We hypothesize that multi-object makes exploring lid-twisting behavior an easier process by introducing an object curriculum that covers both easy and hard object instances during training.
VI-C Qualitative Results
In this section, we study the reward design by examining their induced behavior. Specifically, we compare our method to the reduced contact reward baseline and the gaining reward baseline. Visualization of trajectories from policies trained with each method is shown in Figure 6. Policies trained without the finger contact reward exhibit extremely unstable grasping or screwing motions; the behaviors not only appear unnatural in simulation, but also completely fail to transfer to a real-world setup. We hypothesize that the use of two hands, each with a high degree of freedom, leads to a massive search space for in-hand object manipulation; introducing task priors like behavior-aware contact reward is therefore essential for reducing the search space and eliciting natural behavior.
TABLE II: Comparison with baselines on real setup. For each method, we deploy 3 policies trained on 3 different seeds and average the results. Each deployment trial is conducted for 30 seconds. The Angular Displacement (AD) metric refers to number of degrees that object lid is twisted by hands in a single trial. The Time-to-Fail (TTF) metric refers to number of seconds that has passed before the object drops out of robot hands; this number is only counted when there is such a failure. Our success policy outperforms all baselines on both metrics.
BlueBottle WoodBottle RedBottle GoldBottle SqaureBottle
Method AD (deg) $\uparrow$ TTF (s) $\uparrow$ AD (deg) $\uparrow$ TTF (s) $\uparrow$ AD (deg) $\uparrow$ TTF (s) $\uparrow$ AD (deg) $\uparrow$ TTF (s) $\uparrow$ AD (deg) $\uparrow$ TTF (s) $\uparrow$
Replay 128.33$\pm$217.96 7.67$\pm$4.93 2.67$\pm$4.62 7.67$\pm$5.86 15.00$\pm$25.98 4.67$\pm$4.62 28.33$\pm$43.04 7.67$\pm$4.04 29.67$\pm$8.62 10.00$\pm$0.00
No-Vis 1.33$\pm$2.31 21.67$\pm$14.43 1.07$\pm$1.85 14.67$\pm$13.61 1.90$\pm$3.29 8.33$\pm$6.11 0.67$\pm$1.15 16.33$\pm$13.05 5.00$\pm$6.24 20.33$\pm$11.24
No-Asym 18.67$\pm$28.94 30.00$\pm$0.00 0.67$\pm$1.15 19.33$\pm$15.14 8.33$\pm$14.43 13.00$\pm$15.13 4.3$\pm$7.51 3.67$\pm$3.79 0.00$\pm$0.00 2.33$\pm$1.53
Large 2.00$\pm$2.00 22.33$\pm$13.28 0.00$\pm$0.00 24.00$\pm$10.39 2.33$\pm$2.52 9.33$\pm$4.73 2.67$\pm$3.06 22.33$\pm$13.28 1.67$\pm$2.89 30.00$\pm$0.00
Ours 946.33$\pm$383.81 23.67$\pm$10.97 499.50$\pm$578.23 30.0$\pm$0.0 150.67$\pm$113.47 30.00$\pm$0.00 98.67$\pm$66.91 30.00$\pm$0.00 43.00$\pm$12.12 30.00$\pm$0.00
VII Real-world Experiments
In this section, we first demonstrate that our final learned policy can be directly transferred to the real world in a zero-shot manner. Then, we investigate the effect of several key design choices on the success of sim-to-real transfer. Finally, we examine our policy’s ability to generalize, as well as its robustness against various kinds of perturbations.
VII-A Setup
For quantitative evaluation, we evaluate the sim-to-real transfer capabilities of our policies on five different articulated bottle objects (Figure 4). Among them, four are in-distribution (round-body bottles) and one is outside of the training distribution (square-body bottle). We measure both AD and TTF in 20 trials, with each trial lasting for a maximum of 30 seconds. For each evaluated method, we select the three best policies out of ten policies trained on ten different random seeds. We end a trial if the bottle falls off the palm.
Baselines. We compare our final policy with the following baselines to study the effect several key design choices in our pipeline.
Open-loop Replay Policy (Replay). We record successful trials of our learned policy in the simulation and randomly select a trajectory to replay on the real robot. This baseline is used to evaluate whether the task can be solved by a deterministic motion pattern.
Policy without Vision (No-Vis). This baseline policy only takes proprioceptive state information as input, without information about the object state.
Policy without Asymmetric Training (No-Asym). We compare with a baseline where policy is trained without asymmetric PPO, and evaluate whether introducing additional privielged information into the value network will affect the transfer performance.
Larger Policy Network Size (Large). We increase the size of our actor-network and train a reduced-size policy. We use this to evaluate whether over-parameterization harms policy performance.
VII-B Sim-to-Real Results
We show quantitative results comparing our policy with baseline policies in Table VI-C. For both metrics, our policy outperforms all baselines across all evaluated objects. Our method also achieves a stable grasp and can rotate 3 out of 5 objects at a reasonable speed. In particular, for the blue bottle, one of the deployed policies can achieve 4 full turns (360 degrees) in 30 seconds on average. In contrast, almost all the baselines fail to achieve any effective rotations, either getting stuck or dropping the bottle to the ground. We find that the open-loop policy has the lowest TTF score. Replaying a successful trajectory will not lead to a stable grasp for most of the time, and the bottle will directly roll on the fingers and then drop off the palm. This suggests that the considered task involves very fine-grained contacts and requires the policy to act very precisely according to the object state. Another interesting observation is that the large policy does not transfer to the real world, although we confirm that it can achieve similar performance to our full policy in simulation. This suggests that some overfitting occurs, and controlling the size of the policy network is very important for the successful sim-to-real transfer of our considered contact-rich task.
VII-C Generalization to Novel Objects
We further test our policy’s ability to generalize by testing it on an additional 10 novel objects commonly found in households (Figure LABEL:fig:teaser). These objects substantially differ from our training objects in terms of shape, size, mass, material, and color. Besides, they are also fundamentally different in terms of mechanical design. While the lids of the synthetic bottles that we use for both simulation training and real-world testing can be twisted infinitely, the lids of these household objects cannot. Therefore, to evaluate our policy’s ability to generalize the lid-twisting skill to these novel objects, we use lid removal as a success criterion. We define lid removal as the object lid being completely detached from the object body, as can be seen in the case when the lids fall from the robot hands in Figure LABEL:fig:teaser. We find that our policy can achieve about a 30% success rate. The results can be found in the video supplementary materials.
VII-D Robustness against Perturbation
Finally, we also evaluate our policy’s robustness against force perturbation. Specifically, we perturb the object during deployment at random times by poking or pushing it along random directions using a picker tool (see the left of Figure 7). We find that our policy can reorient and translate the object back to the center for continual manipulation, indicating that it has some robustness to external forces and can adapt to these unexpected changes. Note that we use a marker-based object detection system in this experiment to disentangle the visual occlusion effect.
Figure 7: Perturbing a learned policy with random external force. Our policy is resilient to these external forces and recovers to normal. See our website for video visualizations.
VIII Conclusion
In this paper, we consider the task of twisting or removing lids of bottle-like objects with two hands. We present a RL-based sim-to-real system to solve this problem, and propose various techniques to handle the challenges that arise – including a novel reward design, a sparse object representation for real-time perception, and an efficient yet high-fidelity method to simulate twisting bottle caps. We conduct experiments in both simulation and real world to demonstrate the effectiveness of our approach. Our real-world results show generalization across a wide range of seen and unseen objects. We hope our system can inspire future researchers to tackle bimanual dexterous hands challenges with RL and sim-to-real.

### v2

@@ 5.1 Experiment Setup
Object Set. For quantitative evaluation, we evaluate the sim-to-real transfer capabilities of our policies on five different articulated bottle objects (Figure 2(C)). Among them, four are in-distribution (round-body bottles) and one is outside of the training distribution (square-body bottle).
Evaluation Metric. We measure both AD and TTF in 20 trials, with each trial lasting for a maximum of 30 seconds. For each evaluated method, we select the three best policies out of ten policies trained on ten different random seeds. We end a trial if the bottle falls off the palm.
Baselines. We compare our final policy with the following baselines to study the effect of several key design choices. 1) Open-loop Replay Policy (Replay). We record successful trials of our learned policy in the simulation and randomly select a trajectory to replay on the real robot. This baseline is used to evaluate whether the task can be solved by a deterministic motion pattern. 2) Policy without Vision (No-Vis). This baseline policy only takes proprioceptive state information as input, without information about the object state. 3) Policy without Asymmetric Training (No-Asym). We compare with a baseline where policy is trained without asymmetric PPO, and evaluate whether introducing additional privielged information into the value network will affect the transfer performance. 4) Larger Policy Network Size (Large). We increase the size of our actor-network and train a large-size policy. We use this to evaluate whether over-parameterization harms policy performance.
5.2 Twisting Lids in the Real World
We show quantitative results comparing our policy with baseline policies in Table 4.2. For both metrics, our policy outperforms all baselines across all evaluated objects. Our method can perform stable grasp on all the objects, and can rotate 3 out of 5 objects at a reasonable speed. In particular, for the blue bottle, one of the deployed policies can achieve 4 full turns (360 degrees) in 30 seconds on average. In contrast, almost all the baselines fail to achieve any effective rotations, either getting stuck or dropping the bottle to the ground. We find that the open-loop policy has the lowest TTF score. Replaying a successful trajectory will not lead to a stable grasp for most of the time, and the bottle will directly roll on the fingers and then drop off the palm. This suggests that the considered task involves very fine-grained contacts and requires the policy to act very precisely according to the object state. Another interesting observation is that the large policy does not transfer to the real world, although we confirm that it can achieve similar performance to our full policy in simulation. This suggests that some overfitting occurs, and controlling the size of the policy network is very important for the successful sim-to-real transfer of our considered contact-rich task.
5.3 Robustness against Perturbation
Finally, we also evaluate our policy’s robustness against force perturbation. Specifically, we perturb the object during deployment at random times by poking or pushing it along random directions using a picker tool (see the right of Figure 5). We find that our policy can reorient and move the object back to a stable pose for continuous manipulation, indicating that it has some robustness against external forces and can adapt to these unexpected changes. Note that we use a marker-based object detection system in this experiment to disentangle the visual occlusion effect.
5.4 Exploration of Twisting Lids Off
In the above section, we mainly study whether the twisting behavior can naturally emerge and be transferred to the real world. Next, we explore the limit of our approach by testing it on 10 novel household objects (Figure 2(D)). These objects differ substantially from our training objects in terms of shape, size, mass, material, color, and mechanical design. While the lids of the synthetic bottles that we use for both simulation training and real-world testing can be twisted infinitely, the lids of these household objects cannot. To evaluate our policy’s ability to generalize the lid-twisting skill to these novel objects, we use the success rate on a novel yet adjacent lid-removal task as the criterion. We define lid-removal as the object’s lid being completely detached from the object body (e.g., when the lids fall from the robot’s hands in Figure 1.
We find that our policy continues to achieve stable and natural twisting behaviors on these novel objects. Furthermore, while we only train the policy for lid-twisting, we also find our policy capable of removing lids. For HairMask and FiberGummies, our policy showcases lid-removal rates of more than 50%. For particularly challenging objects that require many turns to remove the lid, such as PeanutButter and EmptyNutella, our policy only achieves 10% lid-removal rates; however, its twisting behavior is steady and robust to perturbation (see our video supplementary for visualization).
6 Conclusion
We present an RL-based sim-to-real system for twisting or removing lids of bottle-like objects with two hands. We propose several techniques to handle the challenges that arise: a novel reward design, a sparse object representation for real-time perception, and an efficient yet high-fidelity method to simulate twisting bottle caps. We conduct experiments in both simulation and real world to demonstrate the effectiveness of our approach. Our real-world results show generalization across a wide range of seen and unseen objects.